# Real-Time Incremental Explanations for Object Detectors

### v1

@@ Abstract
Existing black box explainability tools for object detectors rely on multiple calls to the model, which prevents them from computing explanations in real time. In this paper we introduce incx, an algorithm for real-time incremental approximations of explanations, based on linear transformations of saliency maps. We implement incx on top of d-rise, a state-of-the-art black-box explainability tool for object detectors. We show that incx’s explanations are comparable in quality to those of d-rise, with insertion curves being within 8%, and are computed two orders of magnitude faster that d-rise’s explanations.
@@ 1 Introduction
Figure 1: Detection of an object ‘car’ (Frames) at four different time stamps, with d-rise saliency landscapes and the approximate landscapes generated by incx, showing the similarity between the fully computed d-rise saliency maps and the estimated incx ones. $V_{t}$ represents the object at time $t$.
@@
AI models are now a primary building block of most computer vision systems. However, the inherent opacity of some of these models (e.g., neural networks) together with the sharp increase in their size and complexity, slow their adoption, especially in safety-critical applications, due to the need to understand the reasoning of the model (Ananthaswamy Anil 2023). Explainable AI (XAI) is emerging as a promising solution to mitigate the aforementioned “black box” problem, providing insights into the decision-making process of machine learning models. XAI can be further divided into “white-box” and “black-box” methods. White-box methods have access to the internals of the model, and are hence tailored to a particular model architecture, whereas black-box methods can only access model outputs and are agnostic to the internal implementation of the model.
Image classification is of particular interest in the context of XAI, as images do not, as a rule, have clearly defined sets of features that an AI model uses for their classification. Object detectors are a particular type of image classification tool which detect and classify objects in real-time, making them essential for tasks such as autonomous driving, as well as for video analysis. Many XAI techniques can be adapted to explaining object detectors: in particular, d-rise (Petsiuk et al. 2021) builds upon rise (Petsiuk, Das, and Saenko 2018) by implementing detection vectors that encode the bounding box locations and the classification information. Unfortunately, algorithms used by black-box XAI tools require multiple calls to the model on each input, rendering them unsuitable for real-time applications. Even a state-of-the-art XAI tool such as d-rise takes $0.02$ seconds to produce an explanation of a single frame of video (Kuroki and Yamasaki 2024). Using another model to estimate explanations for the first model, as done by fastshap (Jethani et al. 2022) is even more problematic, using a black box – another model – to estimate the explanation of the first black box. Such an approach leaves the explainer model unexplained.
@@
We introduce incx, or Incremental eXplanations, a black-box XAI algorithm and tool for explaining object detectors in real-time that avoids multiple calls to the model and incurs a negligible overhead over the object classification time. This allows near real-time explanation processing for video data and efficient processing of large amounts of time series image data, making incx useful for explanations of autonomous driving decisions among other applications, for example, an efficient detection of the best locations to place advertisements in video content. incx tracks saliency maps and explanations from one frame to the next, applying translations and scaling to the saliency map generated in the first frame. Unlike fastshap, it does not rely on a separate machine learning model that would necessitate its own explanations.
@@
We show that incx computes explanations that are comparable in quality to d-rise, but outperforms d-rise in speed by two orders of magnitude (see Figure 1 for one example of a comparison between d-rise and incx). To the best of our knowledge, this is the first black-box real-time XAI tool for object detectors.
@@
The code, the benchmark sets, proofs and the full tables of results are submitted as a part of the supplementary material.
@@ 2 Related Work
We survey the related work in object detection, tracking, and explainability. Since this work focuses on black-box explainability, we do not include white-box XAI methods.
@@
There is a large body of work on explainability for image classifiers. Common black-box tools are lime (Ribeiro, Singh, and Guestrin 2016), integrated gradients (Sundararajan, Taly, and Yan 2017), shap (Lundberg and Lee 2017), rise (Petsiuk, Das, and Saenko 2018), and rex (Chockler, Kelly, and Kroening 2023). These tools and algorithms define and output explanations in different formats: functions, models and landscapes. There is a straightforward conversion of different formats of explanations to saliency maps, which rank input pixels based on their contribution to the model classification. Black-box tools generally rely on multiple calls to the model, computing explanations from these results, and hence incurring a significant overhead on top of the image classifier.
Object detection
The object detection task is focused in classifying and finding objects under low latency.
yolo, a collection of one-stage algorithms, has enjoyed a considerable popularity due to its balance between efficiency and accuracy. It divides the image into a grid of cells and assigns a set of class probabilities, a bounding box, and a confidence value. Then it performs non-maximum suppression to remove overlapping bounding boxes. The latest, $10$th version addresses many of the shortcomings of the original yolo (Wang et al. 2024). faster r-cnn (Ren et al. 2015) is a detection algorithm built on r-cnn (Girshick et al. 2014) and fast r-cnn (Girshick 2015). The two-stage process uses a region proposal network (RPN) to identify regions within the image that contain an object. It then takes those regions and warps them into a classifier network to find the probabilities of the objects inside the image.
Transformer-based detectors have been popular since the introduction of the transformer architecture, which revolutionized the Natural Language Processing (NLP) domain, as it introduced the self-attention mechanism (Vaswani et al. 2017). The Real-Time Detection Transformer (rt-detr) leverages the same concept for object detection tasks (Zhao et al. 2024).
@@
Several XAI algorithms can be modified to provide explainability for object detectors. There are only a few black-box algorithms that generate saliency maps without relying on the model architecture. d-rise (Petsiuk et al. 2021) is a modification of rise for black-box object detectors, using the basic underlying idea of rise to generate a set of random masks (Petsiuk, Das, and Saenko 2018). d-rise also introduces the concept of a detection vector, composed of localization information (bounding boxes), object score, and classification information. It then uses this vector to compute a weight for each mask and construct a saliency map given by the weighted sum of masks. In contrast, fsod (Kuroki and Yamasaki 2024) adapts the loss function of fastshap to train an explainer model with an architecture inspired by UNet to account for classification and location of the detected object (Jethani et al. 2022). Although this approach exhibits a considerable faster performance than the previously mentioned ones, its major drawback is that lacks explainability for the explainer model itself. Hence, it is impossible to verify that the generated Shapley values are correct, unless they are computed explicitly or by executing another XAI tool on the explainer model.
@@
Detection-based tracking (DBT) is the dominant paradigm in multiple object tracking tasks. These methods divide the tracking sub-problem into three main tasks: object detection, re-identification (ReID) and trajectory tracking (Du et al. 2024). A classic example of a Multi-object tracker within this paradigm is SORT (Bewley et al. 2016), which is the algorithm we use in this paper. SORT uses the Kalman filter to perform estimations of previously detected objects. Subsequently, it associates the identified objects via the Intersection over Union (IoU) of the bounding box of the detected object and the estimated position and scale of the bounding box. It uses the Hungarian algorithm to optimally solve this assignment. There is a number of algorithms implementing improvements to SORT, such as DeepSORT (Wojke, Bewley, and Paulus 2017), StrongSORT (Du et al. 2023), and ByteTrack (Zhang et al. 2022), but these rely on an additional deep learning model, detracting from transparency.
Interpretable decision making
Interpretable decision-making in autonomous driving has a similar motivation to our work – real-time explanations of AI model’s decisions – but focuses on providing textual explanations to the model’s actions, rather than visual explanations to the images detected by the vehicle’s sensors (Xu et al. 2020; Jing et al. 2022).
@@ 3 Incremental Explanations
In this section we present the theoretical foundations for our saliency map and explanation transformation and the algorithm incx, which employs this transformation to avoid calls to the object detector in subsequent frames. An initial saliency map is assumed to be computed by an external XAI tool. incx is agnostic to the XAI tool as well as to the model, as it does not assume access to the model internals or a specific XAI algorithm. We require only that the XAI output is a saliency map.
@@ 3.1 Theoretical Foundations
The intuition behind our algorithm is based on the following observation. As long as the detected object moves in 3D without turning or deforming, its projection on 2D can only undergo a combination of scaling and linear translation (for instance, if a car gets closer, its 2D projection scales up). We prove that given such a transformation of the 2D projection of the detected object and the initial saliency map, the saliency map of the transformed object can be computed from the original saliency map using a similar transformation, and the initial explanation, mapped into the transformed saliency map should remain valid. The formal proof is based on known results from probability theory and is quite involved (see the supplementary material).
The core of our approach is treating the saliency map as a probability mass function (pmf) and then applying the well-understood mathematics of pmf transformation to estimate a pmf that represents the new, updated, saliency map. First, we need to convert our initial saliency map into a pmf. This is easily done by normalising the values of the saliency map so that they sum to $1$.
Now we use Theorem 1 (Hogg, McKean, and Craig 2019) to help us define the transformations of the pmf. This allows us to calculate the new saliency pmf after undergoing linear transformation (see the supplementary material for details). Technically, our pmf is a joint probability mass function $P(X,Y)$. We restate the core transformation result over a single variable from Hogg, McKean, and Craig (2019) here for completeness.
Theorem 1.
(Hogg, McKean, and Craig 2019) Assume $X$ to be a discrete random variable with space $D_{X}$ and $g$ a one-to-one transformation function. Then the space of $Y$ is $D_{Y}=\{g(x):x\in D_{X}\}$, and the pmf of $Y$ is
@@
Theorem 1 can be extended to joint pmfs over multiple variables (Hogg, McKean, and Craig 2019, p.100) (see supplementary material).
Having established the intuition behind the transformation of the saliency pmf, we address the considerations of our approach concerning transformations involving three-dimensional objects. Specifically, we first describe how such objects get projected onto the image plane. By Forsyth and Ponce (2002), the mapping of points from three-dimensional world coordinates to two-dimensional screen coordinates can be achieved while using a common frame of reference for the world and the camera. Lemma 1 formalizes this projection process.
Lemma 1 (3D world coordinates to 2D screen (camera image) coordinates).
@@
Furthermore, we formalize the center of the bounding box for the projected object in Definition 1. This concept is leveraged in the scaling and translation operations described in Definition 2, which are fundamental for incx.
@@
Let $A=\{(u,v)\in\mathbb{R}^{2}\}$ represent a set of points defining any area in the image space.
Scaling is a linear transformation $S:A\to\mathbb{R}^{2}$, defined as $S(\mathbf{p})=\gamma(\mathbf{p}-\mathbf{C}(A))$, where $\mathbf{p}\in A$ and $\gamma\in\mathbb{R}^{2\times 2}$ is a diagonal matrix.
Translation is a linear transformation $T:A\to\mathbb{R}^{2}$, defined as $T(\mathbf{p})=\mathbf{p}+\mu$, where $\mathbf{p}\in A$.
The scaling transformation $S$ scales $\mathbf{p}$ by a factor $\gamma$ after translating the center $\mathbf{C}(A)$ to the origin $(0,0)$. The translation function $T$ shifts the two-dimensional input $p$ by a vector $\mu\in\mathbb{R}^{2}$.
[…]
Notation: Let $\mathbf{Q}_{t}=(Q_{1,t},Q_{2,t})$ be a vector in the image space at time $t$, where $Q_{1,t}$ and $Q_{2,t}$ represent random variables along the $u$-axis and the $v$-axis, respectively. The pmf of $\mathbf{Q}_{t}$ is $p_{\mathbf{Q}_{t}}(\mathbf{q}_{t})$, which is derived from the normalized saliency map.
@@
where $\mathbf{t}\in\mathbb{R}^{3}$ is an arbitrary translation vector, $\mathbf{I}\in\mathbb{R}^{3\times 3}$ is the identity matrix, and $m_{x}$, $m_{y}$, and $m_{z}\in\mathbb{R}$ are scaling factors.
Finally, let $S$ be the scaling function, $T$ the translation function, $\mathbf{q}_{t+1}$ a coordinate vector in the image space at time $t+1$, and $\mathbf{q}_{t}$ a coordinate vector at time $t$. The following theorem describes the relation between the pmf at time $t$ and at time $t+1$ and is instrumental to the incx algorithm.
@@ 3.2 Algorithm Structure
The incx algorithm is schematically presented in Figure 2. We list three object detectors, but, as we mentioned earlier, incx is agnostic to the detector, as it treats it as a black box. As previously stated, the goal of our algorithm is to approximate saliency maps and explanations in real-time for object detection tasks. As a result, we need a starting point from where we can move forward. Hence, the saliency map in the first frame is computed using an XAI tool (Figure 2 lists d-rise, but it can be any saliency map-producing tool). After this bootstrapping, the goal is to identify a way to extend the saliency map obtained in the first frame and extrapolate from it to create subsequent ones. To achieve this, we perform the transformation defined in Theorem 2.
The first involves bootstrapping our method by computing the initial saliency map. This process is quite straightforward and involves calling an object detector $\mathcal{N}$ and explainer $h$. Only the saliency maps found in the first frame will be used as a basis for extrapolation in later frames. For the subsequent frames, we need to implement a different approach to be able to obtain approximate saliency maps in real-time. First we incorporate a new block (see Figure 2), which is the object tracker: SORT (Bewley et al. 2016). This object tracker takes the bounding box of the identified objects and checks if the id corresponds to any of the objects identified in the previous frame (Line 11). Once an object has been identified, the linear transformation is performed (Lines 1216). This process takes the previous bounding box and the current bounding box and computes $\Delta x$, $\Delta y$, $\gamma$. These computed transformations are applied to the saliency map, utilizing  Theorem 2. We note that when the saliency map is scaled up, gaps may appear in the newly computed pmf, requiring interpolation to fill them.
@@
It is important to note that we can track explanations only as long as the object tracker finds the objects identified in the first frame. An object may vanish from the frame or remain undetected for a long time. For reasons of efficiency, we define a timeout that allows us to stop attempting to explain an object which is no longer detectable (Line 20).
Algorithm 1 Incremental Explanation
@@
Input: A video $v$, an object detector $\mathcal{N}$, an explainer $h$
Output: A sequence of saliency maps $s$ and explanations $xp$
1:  $x_{t-1},y_{t-1},\text{width}_{t-1},\text{height}_{t-1}\leftarrow 0,0,0,0$
2:  $\text{id}_{0},\text{count}\leftarrow 0,0$
3:  initialize empty saliency map $s_{t-1}$
4:  initialize empty arrays $s$, $xp$
5:  for each frame $f_{t}$ at time $t$ in video $v$ do
6:     $(x_{t},y_{t},\text{width}_{t},\text{height}_{t})\leftarrow\mathcal{N}(f_{t})$
7:     if $t==0$ then
8:        $s_{t}\leftarrow h(f_{t})$
9:     else
10:        $\text{id}_{t}\leftarrow SORT(f_{t})$
11:        if $\text{id}_{0}==\text{id}_{t}$ then
12:           $\Delta u\leftarrow u_{t}-u_{t-1}$
13:           $\Delta v\leftarrow v_{t}-v_{t-1}$
14:           $\gamma\leftarrow(\frac{\text{width}_{t}}{\text{width}_{t-1}},\frac{\text{height}_{t}}{\text{height}_{t-1}})$
15:           $s_{t}\leftarrow\text{Scale}(s_{t},\gamma)$
16:           $s_{t}\leftarrow\text{Translate}(s_{t-1},\Delta u,\Delta v)$
17:           $\text{count}\leftarrow 0$
18:        else
19:           $\text{count}\leftarrow\text{count}+1$
20:           if $\text{count}>=\emph{timeout}$ then
21:              Restart
22:           end if
23:        end if
24:     end if
25:     $xp_{t}\leftarrow\text{Explanation}(s_{t},\mathcal{N},f_{t},t)$
26:     $xp\leftarrow xp+[xp_{t}]$
27:     $s\leftarrow s+[s_{t}]$
28:     $\text{width}_{t-1},\text{height}_{t-1}\leftarrow\text{width}_{t},\text{height}_{t}$
29:     $x_{t-1},y_{t-1}\leftarrow x_{t},y_{t}$
30:     $s_{t-1}\leftarrow s_{t}$
31:  end for
32:  return $s,xp$

### v2

@@ Abstract
Object detectors are widely used in safety-critical real-time applications such as autonomous driving. Explainability is especially important for safety-critical applications, and due to the variety of object detectors and their often proprietary nature, black-box explainability tools are needed. However, existing black-box explainability tools for AI models rely on multiple model calls, rendering them impractical for real-time use.
In this paper, we introduce IncX, an algorithm and a tool for real-time black-box explainability for object detectors. The algorithm is based on linear transformations of saliency maps, producing sufficient explanations. We evaluate our implementation on four widely used video datasets of autonomous driving and demonstrate that IncX’s explanations are comparable in quality to the state-of-the-art and are computed two orders of magnitude faster than the state-of-the-art, making them usable in real time.
@@ 1 Introduction
Figure 1: Detection of an object ‘car’ (Frames) at four different time stamps, with d-rise saliency landscapes and the approximate landscapes and explanations generated by IncX, showing the similarity between the fully computed d-rise saliency maps and the estimated IncX ones. $V_{t}$ represents the object at time $t$.
@@
AI models are now a primary building block of most computer vision systems. However, the inherent opacity of some of these models (e.g., neural networks) and their sharp increase in size and complexity, slow their adoption, especially in safety-critical applications, due to the need to understand the reasoning of the model [2].
Explainable AI (XAI) is used to provide insights into the decision-making process of machine learning models. Post hoc XAI can be roughly divided into “white-box” and “black-box” methods. White-box methods have access to the internals of the model, and are hence tailored to a particular model architecture. Black-box methods can only access model outputs and are agnostic to the internal implementation of the model. In this paper we focus on black-box explainability, due to its portability across models.
Explainable AI is particularly relevant in the context of autonomous driving, as it is instrumental in assessing safety of autonomous vehicles. Regulatory bodies can greatly benefit from XAI by improving model explainability, which aids in product liability litigation and compliance assessments [10]. Furthermore, industry standards for road vehicle safety, such as ISO21448 [21], emphasize the need for traceability of performance insufficiencies and their trigger conditions. In this regard, explainable AI has been proposed as a means of enhancing implementation transparency, contributing to the development of inherently safe AI systems in autonomous driving (AD) [31].
Object detection is crucial in AD for various tasks, including traffic sign and light recognition [24]. Visual explanations, such as saliency maps, are essential for enhancing transparency and trust in autonomous driving systems [3]. They aid in generating textual observations and improving system interpretability [25]. Explaining object detectors, however, is challenging. On the one hand, most AD approaches rely on white-box algorithms [33, 5], which are primarily based on gradient propagation, restricting their applicability to specific architectures. On the other hand, while black-box methods are model-agnostic, their high computational cost makes real-time explanations infeasible [32].
@@
To enhance autonomous driving models’ explainability, we introduce Incremental eXplanations, or IncX —a black-box XAI algorithm and tool for explaining object detectors in real-time. IncX avoids multiple calls to the model and thus has a negligible overhead over the object classification running time, allowing near real-time explanation processing for video data. IncX tracks saliency maps from one frame to the next, applying linear transformations to the saliency map generated in the first frame, and uses these maps to compute sufficient explanations—subsets of pixels of the image sufficient for the original classification. Unlike fastshap [22], it does not rely on a separate machine learning model that would necessitate, in turn, its own explanations.
@@
We show that IncX computes explanations that are comparable in quality to d-rise [36], but outperforms d-rise in speed by two orders of magnitude. To the best of our knowledge, this is the first black-box real-time XAI tool for object detectors.
@@
The code, the benchmark sets, full proofs, and the full tables of results are submitted as a part of the supplementary material. Instructions for running IncX locally in a docker container and seeing the explanations in real time are also provided.
@@ 2 Related Work
We survey the related work in object detection, tracking, and explainability. Since this work focuses on black-box explainability, we do not include white-box XAI methods, such as GradCAM [41] or DexT [34].
@@
There is a large body of work on explainability for image classifiers. Common black-box tools are lime [39], shap [30], rise [35], and rex [9]. These tools and algorithms define and output explanations in different formats: functions, models and landscapes. There is a straightforward conversion of different formats of explanations to saliency maps, which rank input pixels based on their contribution to the model classification. Black-box tools generally rely on multiple calls to the model, computing explanations from these results, and hence incurring a significant overhead on top of the image classifier.
Object detection in AD
The object detection task is finding objects and classifying them under low latency. This capability is fundamental in autonomous driving, as it enables the vehicle to recognize other automobiles and traffic signs in real time.
yolo [37], a collection of one-stage algorithms, has enjoyed a considerable popularity due to its balance between efficiency and accuracy, leading to adoption in autonomous driving research [1]. It divides the image into a grid of cells and assigns a set of class probabilities, a bounding box, and a confidence value. Then it performs non-maximum suppression to remove overlapping bounding boxes. The latest version, $10$th as of time of writing, addresses many of the shortcomings of the original yolo [43].
In contrast, a two-stage process uses a region proposal network (RPN) to identify regions within the image that contain an object. Then it takes those regions and warps them into a classifier network to find the probabilities of the objects inside the image. Examples of such architectures include faster r-cnn [38], r-cnn [18] and fast r-cnn [17]. This approach has also been extensively used in AD research [28].
Transformer-based detectors have been popular since the introduction of the transformer architecture, which revolutionized the Natural Language Processing (NLP) domain, as it introduced the self-attention mechanism [42]. The Real-Time Detection Transformer (rt-detr) leverages the same concept for object detection tasks [49].
@@
Several XAI algorithms can be modified to provide explainability for object detectors. There are only a few black-box algorithms that generate saliency maps without relying on the model architecture. d-rise [36] is a modification of rise for black-box object detectors, using the basic underlying idea of rise to generate a set of random masks. d-rise also introduces the concept of a detection vector, composed of localization information (bounding boxes), object score, and classification information. It then uses this vector to compute a weight for each mask and construct a saliency map given by the weighted sum of masks. In contrast, fsod [26] adapts the loss function of fastshap to train an explainer model with an architecture inspired by UNet to account for classification and location of the detected object [23]. Although this approach exhibits a considerably faster performance than the previously mentioned ones, its major drawback is that it lacks explainability for the explainer model itself.
@@
Detection-based tracking (DBT) is the dominant paradigm in multiple object tracking tasks. These methods divide the tracking sub-problem into three main tasks: object detection, re-identification (ReID) and trajectory tracking [12]. A classic example of a Multi-object tracker within this paradigm is SORT [4], which is the algorithm we use in this paper. SORT uses a Kalman filter to perform estimations of previously detected objects. Subsequently, it associates the identified objects via the Intersection over Union (IoU) of the bounding box of the detected object and the estimated position and scale of the bounding box. It uses the Hungarian algorithm to optimally solve this assignment. There is a number of algorithms implementing improvements to SORT, such as DeepSORT [46], StrongSORT [13], and ByteTrack [48], but these rely on an additional deep learning model, detracting from transparency.
Explainable AI in Autonomous Driving
Explainable AI for safe autonomous driving can be categorized into several approaches [27]. Interpretable by design models are inherently transparent, such as traffic sign detectors built with Inductive Logic Programming [7]. Interpretable surrogate models approximate black-box behavior with explainable alternatives like LIME [39]. Interpretable monitoring employs saliency maps to track model behavior and detect anomalies [19]. Interpretable validation ensures reliability through offline interpretable methods. Together, these techniques strengthen AI transparency and safety in autonomous driving.
@@ 3 Incremental Explanations
In this section we present our contribution – the IncX algorithm. Full proofs of all theorems and lemmas are provided in the supplementary material.
@@ 3.1 Theoretical Foundations
The intuition behind our algorithm is based on the following lemma.
Lemma 1.
For a fixed observer, movement of an object in $3D$ space without rotation or deformation can only result in a combination of scaling and linear translation when projected on a given vertical plane.
Proof Sketch.
The lemma follows from the observation that the vector of movement of an observed object is a sum of its projections to the vertical plane and in the direction to or away from the observer. The vector to or away from the observer determines the scaling, while the projection to the vertical plane determines the linear translation. ∎
Given a transformation of the observed object as above and its initial saliency map, we prove that the saliency map of the transformed object can be computed from the original saliency map using a similar transformation.
The main result is stated in Theorem 1. The formal proof is based on results from probability theory and is quite involved (see the supplementary material). Before presenting the theorem, we first introduce additional context and formalization to ensure a clearer understanding.
Intuitively, the proof is based on converting the saliency map to a probability mass function (pmf) and then applying the mathematics of pmf transformation to estimate a pmf for a new, updated, saliency map.
First, the initial saliency map is converted into a pmf by normalizing the values of the saliency map. Then we use  Lemma 2 [20] to define the transformations of the pmf. This allows us to calculate the new saliency pmf after undergoing linear transformation. Technically, our pmf is a joint probability mass function $P(X,Y)$. We restate the core transformation result over a single variable from Hogg et al. [20] here for completeness.
Lemma 2 ([20]).
Assume $X$ to be a discrete random variable with space $D_{X}$ and $g$ a one-to-one transformation function. Then the space of $Y$ is $D_{Y}=\{g(x):x\in D_{X}\}$, and the pmf of $Y$ is
@@
Lemma 2 can be extended to joint pmfs over multiple variables [20, p.100] (see supplementary material).
Having established the intuition behind the transformation of the saliency pmf, we address the considerations of our approach concerning transformations involving three-dimensional objects. Specifically, we first describe how such objects get projected onto the image plane. By Forsyth and Ponce [15], the mapping of points from three-dimensional world coordinates to two-dimensional screen coordinates can be achieved while using a common frame of reference for the world and the camera. Lemma 3 formalizes this projection process.
Lemma 3.
@@
We formalize the intuition of a center of a bounding box for the projected object in Definition 1. This concept is leveraged in the scaling and translation operations described in Definition 2, which are fundamental for IncX.
@@
Let $A=\{(u,v)\in\mathbb{R}^{2}\}$ represent a set of points defining an area in the image space.
Scaling is a linear transformation $S:\mathbb{R}^{2}\to\mathbb{R}^{2}$, defined as $S(\mathbf{p})=\gamma(\mathbf{p}-\mathbf{C}(A))$, where $\mathbf{p}\in A$ and $\gamma\in\mathbb{R}^{2\times 2}$ is a diagonal matrix.
Translation is a linear transformation $T:\mathbb{R}^{2}\to\mathbb{R}^{2}$, defined as $T(\mathbf{p})=\mathbf{p}+\mu$, where $\mathbf{p}\in A$.
The scaling transformation $S$ modifies a point $\mathbf{p}$ by a factor of $\gamma$, after first shifting it relative to the center of the area of the object in the image. Specifically, applying $S$ to all points in $A$ is equivalent to scaling the entire image while moving the center of mass $\mathbf{C}(A)$ to the origin $(0,0)$. The translation function $T$ shifts the two-dimensional point $\mathbf{p}$ by a vector $\mu\in\mathbb{R}^{2}$. Similarly, applying this transformation to all points in an image results in translating the entire image by $\mu$.
We first provide the intuition behind our theorem and then its formal statement in Theorem 1. Our approach maps a given pmf to a $2D$-projection of a three-dimensional object $\delta_{t}$ (see Assumption 1). Furthermore, as long as the object being tracked does not leave the frame, rotate, or is occluded by another object, the future positions of the object are determined solely by translation and scaling transformations in the three-dimensional space (see Assumption 2). Together, these assumptions lead to the statement that future saliency maps can be computed by applying a scaling and translation linear transformation to the current saliency map.
We use the following notation. Let $\mathbf{Q}_{t}=(Q_{1,t},Q_{2,t})$ be a vector in the image space at time $t$, where $Q_{1,t}$ and $Q_{2,t}$ represent random variables along the $u$-axis and the $v$-axis, respectively. The pmf of $\mathbf{Q}_{t}$ is $p_{\mathbf{Q}_{t}}(\mathbf{q}_{t})$, which is derived from the normalized saliency map.
@@
where $\mathbf{t}\in\mathbb{R}^{3}$ is a translation vector, $\mathbf{I}\in\mathbb{R}^{3\times 3}$ is the identity matrix, and $m_{x}$, $m_{y}$, and $m_{z}\in\mathbb{R}$ are scaling factors.
Finally, let $S$ be the scaling function, $T$ the translation function, $\mathbf{q}_{t+1}$ a coordinate vector in the image space at time $t+1$, and $\mathbf{q}_{t}$ a coordinate vector at time $t$. The following theorem describes the relation between the pmf at time $t$ and at time $t+1$ and is instrumental to our algorithm.
@@ 3.2 Algorithm Structure
The IncX algorithm is schematically presented in Figure 2, and its pseudo-code is shown in Algorithm 1. In Figure 2, we list the object detectors used for the experiments in the paper, however IncX is agnostic to the detector, hence any object detector can be plugged in.
The saliency map in the first frame is computed using an XAI tool (in our experiments we used d-rise). The Explanation procedure is used to extract a sufficient explanation from the saliency map (see Section 3.3 below). After this bootstrapping, IncX computes saliency maps and sufficient explanations for subsequent frames by performing the transformation described in Theorem 1 on the output of the previous frames.
As shown in Algorithm 1, in subsequent frames IncX maps the current objects to the objects that are tracked in the previous frame using SORT [4]. Once the mapping is computed, the algorithm computes the linear transformation of the bounding boxes of the mapped objects by calling the Transform function (Line 8). This transformation is applied to the saliency map of the previous frame, using Theorem 1. Note that if the new saliency map is scaled up, the gaps in the new pmf are filled using interpolation.
@@
The algorithm can compute explanations only for objects detected by the object tracker. An object might leave the frame or remain undetected for a period of time and reappear later. For efficiency, we define a timeout that allows us to stop attempting to explain an object, which is no longer detectable.
Algorithm 1 Pseudo-code of IncX
@@
Input: Video $v$, object detector $\mathcal{N}$, explainer $h$, tracker SORT
Output: List of explanations $xp$ and estimated saliency maps $s$
1:  initialize bounding boxes and saliency map
2:  initialize empty saliency list $s$ and explanation list $xp$
3:  for each frame $f_{t}$ at time $t$ in $v$ do
4:     $(x_{t},y_{t},w_{t},h_{t})\leftarrow\mathcal{N}(f_{t})$
5:     if $t=0$ then
6:        $s_{t}\leftarrow h(f_{t})$  {Initial saliency map}
7:     else
8:        $s_{t}\leftarrow\text{Transform}(s_{t-1})${track object(s) from previous frame}
9:     end if
10:     $xp_{t}\leftarrow\text{Explanation}(s_{t},\mathcal{N},f_{t},t)$
11:     $xp\leftarrow xy\cup xp_{t}$
12:     $s\leftarrow s\cup s_{t}$
13:  end for
14:  return $xp,s$